Lithium and nitrogen react in a combination reaction to produce lithium nitride:

6Li (s) + N2 (g) → 2Li3N (s)

In a particular experiment, 2.50-g samples of each reagent are reacted. The theoretical yield of lithium nitride is ________ g.

Question 30 options:

A)

12.5

B)

4.18

C)

2.51

D)

6.2

E)

2.09

Question 31 (2.5 points)
Answer #1

solution

given :

mass of li(s) = 2.5 gram

mass of N2(g) = 2.5 gram

molar mass of lithium = 7 gram/mole

molar mass of N2 = 28 gram/ mole

no of mole of li = given mass / molar mass

no of mole of lithium = 2.5 /7 = 0.36 mole

no of mole of N2 = 2.5 / 28 = 0.09 mole

6 mole of lithium is required for 1 mole of N2 gas

for 1 mole of lithium 1/6 mole of N2 is required

for 0.36 mole of lithium (1/6 )* 0.36 = 0.06 mole of N2

so lithium is limiting reagent.

6 mole of lithium is required for 2 mole of Li3N

for 0.36 mole of lithium (2/6)* 0.36 mole of Li3N will form

mole of Li3N = 0.12 mole

molar mass of Li3N = 35 gram/mole.

so mass of Li3N = no of mole* molar mass

mass of Li3N = 0.12* 35 = 4.2 gram

so theoretical yield of Li3N = 4.2 gram

hope your problem has been solved

thank you